f(x)= 5x-3, g(x)=absolute value of x+2. what is the answer for g[f(0)] and f[g(0)]
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

f(x)= 5x-3, g(x)= |x+2|
what is the answer for:
g[f(0)] = g[-3] = |-3+2| = |-1| = 1
----
f[g(0)] = f|0+2| = f(2) = 5*2-3 = 7
=======================================
Cheers,
Stan H.
===========

Answer by edjones(8007)

f(x)= 5x-3, g(x)=|x+2|
.
g(f(0))=|5x-3+2|=|-3+2|=1
.
f(g(0))=5|x+2|-3=5|2|-3=10-3=7
.
Ed